Make New Entry,
Make Followup Entry
User name M. Cummings
Log entry time 08:08:04 on March 8,2012
Entry number 364548
keyword=Owl Shift Summary
Owl Shift Summary
Shift Workers: Melissa Cummings (SL), Toby Badman (TO), Marco Pannunzio Carmignotto (3rd)
00:40 Beam is back, continue with harp scans.
01:38 Hall C changed half wave plate
02:48 Hall C changed half wave plate
03:30 Pengjia finished with harp scan. Moving on to BPM and slow raster calibration. Turned beam off to move target. Septum is set to 718.1A on left and right.
04:00 Hall C changed half wave plate
04:50 Hall C changed half wave plate
05:00 Still working on BPM/slow raster calibration. Adjusting target position to try and identify beam position.
05:55 Hall C changed half wave plate
06:20 LHRS momentum changed to 2.0934 to try and move away from elastic peak
06:31 Beam down while fixing an RF issue
07:15 Beam is back, continuing with BPM/slow raster calibration
07:57 Hall C changed half wave plate
08:00 Cycling LHRS dipole back to 2.251 in preparation for optics